# Predicted High-Pressure Hot Superconductivity in Li2CaH16 and Li2CaH17 Phases that Resemble the Type-II Clathrate Structure

### v1

@@ 3.2 Ternary Phase Diagrams
The most stable structures found in the CSP searches were employed to construct 3D convex hulls (Figures S1-S3) at 100, 200 and 300 GPa at zero-temperature and within the clamped nuclei approximation. Because the compositions considered all possessed a 2:1 lithium to calcium ratio while the hydrogen content varied, 2D cuts of the generated convex hulls, corresponding to a plane of increasing H content from Li2Ca at 100 GPa, or 2 Li + Ca at 200 and 300 GPa were also plotted (Figure S5, S6). At 300 GPa a Li2CaH17 phase was predicted to be thermodynamically stable, while the structure closest to, but not on, the hull, Li2CaH16, lay 7 meV/atom above it (Table S2 and Figure 1). Notably, as discussed above, both of these possessed hydrogenic clathrate cages.
@@
The zero-point-energy (ZPE) is often important in determining relative energy orderings for systems containing light elements, with the greatest potential influence on H-bearing phases, as they can have the highest frequency vibrations. For a given pressure the ZPE is affected by symmetry, with lower symmetry systems generally having larger ZPEs, and the nature of the hydrogenic motifs present within the crystal, with dihydrogen molecules having a larger ZPE than hydrogens within extended lattices such as clathrate cages. Indeed, it has been shown that the ZPE can affect the identity of the on-hull structures in high-pressure polyhydrides containing electropositive metal atoms 70. Obtaining the ZPE necessitates performing phonon calculations, which allows for confirmation of dynamic stability, indicated by a lack of imaginary modes.
@@
Phonon calculations, performed in the harmonic approximation, suggested that Li2CaH17, with a hydrogenic clathrate lattice, and Li2CaHn, = (10,12,15,18-20), which did not feature clathrate structures, were dynamically stable at 300 GPa. Li2CaH16, on the other hand, possessed two modes with imaginary frequencies. These imaginary frequencies are related to the acoustic modes, with the majority of the atomic displacement being due to hydrogen. These frequencies were found to become real and positive by 350 GPa (Figure S22), indicating stabilization due to pressure. A first-principles molecular dynamics simulation at 330 K, however, showed that the hydrogenic lattice remained intact, with no molecular H units forming. This suggests one or more of the following: the phase is dynamically stable only at finite temperature; the system may be stabilized by the inclusion of anharmonicity or quantum effects; or the fineness of the grid needed to obtain real phonons is prohibitively computationally expensive. Due to the large size of the unit cell of Li2CaH16, and the associated computational expense, further calculations investigating this discrepancy and determining the lower boundary of dynamic stability were not performed, and the ZPE was estimated neglecting the imaginary modes. Figure 1 (bottom panel) illustrates that inclusion of the ZPE lowered the enthalpy of Li2CaH16 so that it joined Li2CaH17 on the 300 GPa hull, but the remaining non-clathrate structures became destabilized, lying at least 40 meV/atom above the hull.
@@
Decreasing pressure moves all of the ternary hydrides progressively further from the hull. Li2CaH16 shifts up to be 16 and 35 meV/atom above it, while Li2CaH17 is calculated to be 9 and 36 meV/atom shy of thermodynamic stability at 200 and 100 GPa, respectively. However, these compounds may still be promising at sub-300 GPa pressures, despite not laying on the hull, provided they are dynamically and thermally stable. Analysis of the zero-temperature DFT-computed enthalpies of the compounds contained in the Materials Project Database led to the estimate that 50% of known inorganic crystalline materials are metastable, with a median metastability energy of 15 meV/atom and a 90 percentile of 67 meV/atom 71. For the specific case of high-pressure hydrides, whose synthesis at high-pressures in diamond anvil cells is often catalyzed by non-equilibrium processes such as thousands-of-Kelvin degree laser heating, numerous examples of synthesized DFT-metastable compounds are known. Some examples include Ca2H567, which lays 20 meV/atom above the hull at 25 GPa, and various PHn phases, which lay 30 meV/atom above the hull within the pressure range 80-225 GPa 72, 73, 74, 75.
@@ 3.3 Hydrogen Clathrate Structures
Plots of the enthalpy versus pressure of the low lying structures found in our CSP searches (Figures S8, S9) showed that Li2CaH16 became preferred (over other phases with this stoichiometry) above 270 GPa, while for Li2CaH17 the structure was the most stable above 140 GPa. From the other candidates found in our EA searches, these had the lowest enthalpies up to a maximum investigated pressure of 500 GPa. The propensity for high in compounds with hydrogenic clathrate lattices, the observation that Li2CaH16 and Li2CaH17 both possessed such lattices with a high degree of symmetry, and the finding that both lay on the 300 GPa convex hull upon inclusion of the ZPE prompted us to examine them in more detail.
@@
Li2CaH16 is isotypic with previously predicted Li2MgH16 and Li2NaH16 41, 43. In it, each Ca atom lies in a cage surrounded by four hexagons and twelve pentagons, made of 28 hydrogen atoms, and each Li atom is enclathrated in an 18 atom hydrogen cage with four hexagons and five pentagons. The size of the cages around each metal atom are commensurate with their radii, with Ca nestled in the larger cage and lithium sitting in the smaller cage, as illustrated in Figure 2. At 350 GPa these polygons are highly regular, with H-H bond lengths of 1.05, 1.08, and 1.12 Å. This symmetry lattice is markedly similar to the type-II SiCL structure, differing by a single Wyckoff position. While the clearest hydrogen motifs in Li2CaH16 are the H18 and H28 cages, we can also view the structure as possessing eight H4 tetrahedra in the conventional cell ( = 8), corresponding to one H4 tetrahedra per formula unit. These H4 tetrahedra are centred around the missing Wyckoff position that if filled would make them isotypic to the type-II SiCL structure.
@@
In Li2CaH17, also shown in Figure 2, each Ca atom is surrounded by four hexagons and twelve pentagons, and each Li atom is once again found in the smaller cage, surrounded by eleven pentagons. The additional hydrogen atom in Li2CaH17 renders the phase nearly isostructural with the type-II SiCLs. In comparison with Li2CaH16, Li2CaH17 possesses H4 tetrahedra “stuffed” with the extra hydrogen atom, which assists in completing the cage around the Li ions. The stuffed hydrogen atom may stabilize this structure by balancing the chemical pressure interactions within the lattice, as has been observed in the H-stuffed H4 tetrahedra of LaH10 60, or by allowing the structure to distort more readily. These distortions can be seen with the reduced symmetry of the faces comprising the polygons that make up the clathrate; each pentagon has H-H bond lengths between 1.08 and 1.14 Å, while each hexagon has H-H bond lengths between 1.02 and 1.14 Å at 300 GPa. Investigation of the Li2CaH17 phase shows that these bonds become less uniform as pressure decreases. The phase was determined to be dynamically stable down to 160 GPa, however the full clathrate structure does not persist to this pressure, due to the H-H bond lengths increasing (e.g. see ELF plot in Figure S21).
@@
Li2CaH17 was not anticipated to be the most stable spacegroup for this stoichiometry. Previous studies have predicted an M3H17 phase, where M is a metal atom, in compounds such as Li2(Sc/La/Y/Na)H17 27, 41. These symmetry M3H17 phases are isostructural with the type-II SiCL, while the Li2CaH17 phase can be seen as a distortion of this structure. The external pressure, internal (or chemical pressure) and the number of valence electrons that are transferred from the metal atoms to the hydrogen cage, and potential metal-hydrogen interactions may be important factors driving this distortion. The external pressure most directly affects the (pressure-volume) term to the enthalpy, often favoring the formation of closed packed structures at higher pressures. The chemical pressure is related to the ionic radius of the metal atom and how it fits into the cage, with ions that are too-small or too-large resulting in distorted lattices 60. The number of valence electrons that are donated from the electropositive element to the hydrogenic lattice can influence the types of hydrogenic motifs that are present 10, and the resulting H-H distances, which are also influenced by interactions such as Ca d H2 backbonding 67. To confirm that the phase is indeed favored over , the cubic structure was built and optimized. Comparison of the 300 GPa enthalpies revealed that the phase possessed a lower internal energy, , but a higher term. The bonds in the phase are more distorted than the highly ordered phase, and these distortions likely contribute to the enthalpic favorability of the structure for the metal atoms studied here. Furthermore, the phase of Li2CaH17 was determined to be both dynamically unstable at 300 GPa, and enthalpically disfavored over the pressure range of 100-500 GPa, so it was therefore not considered for additional calculations.
@@
Compared to molecular H-H bonds at the same pressure, the H-H bond lengths in both Li2CaH17 and Li2CaH16 are longer (Table 1). Bader analysis showed that the caged Li and Ca atoms donate charge to the nearby hydrogen atoms, which may covalently bond with one another. For Li2CaH17 at 160 GPa, each H atom gains on average 0.16, with Ca losing 1.05 and Li donating a smaller amount of charge, 0.84. When the pressure is increased to 300 GPa, each H atom in Li2CaH17 gains on average 0.15, while Ca loses 0.89 and Li donates 0.80. Thus, increasing pressure appears to even out the amount of charge transferred from both the formally divalent and monovalent metal atoms, perhaps because a shorter distance between the hydrogen atom and calcium can enhance the amount of backdonation between the occupied antibonding hydrogen-lattice states and the Ca d orbitals, as previously found within compressed CaH4 67 and Ca-S-H ternary hydrides 76. The Bader charges calculated for Li2CaH16 at 350 GPa are similar, with H gaining 0.15 on average, and Ca and Li donating 0.90 and 0.78, respectively.
@@ 3.4 Electronic Structure and Superconductivity
Li2CaH16 and Li2CaH17 are calculated to have a high (Figure 3), primarily of hydrogen character, boding well for superconductivity. The contribution of the calcium states to is intermediate, while the contribution of the lithium states is minimal. Therefore, the DOS plots are consistent with a Li+ cation, and a not-fully-ionised calcium, resulting from the backdonation of charge from the hydrogenic lattice to the Ca d orbitals, as supported by the large amount of Ca d states at . In Li2CaH16 at 350 GPa and Li2CaH17 at 300 GPa lies just below a peak in the DOS, suggesting that further electron doping (e.g. by replacing Li via a divalent metal atom with a similar ionic radius, or Ca by a trivalent metal atom) could potentially enhance . Such peaks in the DOS have previously been reported in symmetry Li2NaH17 (just above ) and Li2YH17 (just below ) at 300 GPa 27, 41. Assuming a full transfer of electrons from the electropositive element to the hydrogenic lattice yields three-donated-electrons in Li2NaH17 and four-donated-electrons in Li2CaH17. In both of these systems sits below the peak in the DOS, so more electrons must be added. However, in Li2YH17, with five-donated-electrons, lies just above the peak. Thus, the optimal number of electron required to maximize is between four and five. We find that adding an additional 0.38 per FU (formula unit) (4 /10 FU) to Li2CaH16 at 350 GPa and an additional 0.40 per FU (4 / 10 FU) to Li2CaH17 at 300 GPa would place on that peak. Therefore, a quaternary compound with an approximate Li2Ca0.6M0.4H16/17 stoichiometry, where M is a trivalent metal ion such as Sc, Y, or La may potentially result in a higher than those calculated here, provided the structure is dynamically stable.
@@
The descriptors for high- superconductivity – and the logarithmic average phonon frequency, – were computed for these two phases at pressures where they were dynamically stable (Table 2). For Li2CaH16, imaginary modes were absent only at 350 GPa, where this phase possessed the largest that we computed. The large of 2.58 necessitated that the Eliashberg equations be solved numerically, and using a value of 0.1 for the Coulomb shielding parameter, , we obtained a of 330 K for Li2CaH16 at this pressure. Though this predicted is over 100 K lower than that computed for the isotypic Li2MgH16 at 250 GPa 43, it surpasses room temperature, rendering Li2CaH16 a “hot” superconductor. Its phonon band structure, phonon projected densities of states and Eliashberg spectral function are shown in Figure 4(a). Because of the mass differences between the three atom types, the phonons can be well separated into regions where the contributions from a particular atom are dominant. Therefore, the total EPC could be partitioned into contributions originating from vibrations in the low-frequency regime, projected mostly onto the calcium atoms (15%), followed by a lithium-dominant region (9%), then by the mainly hydrogenic modes above 750 cm-1 (76%). No high-frequency modes indicative of the presence of molecular hydrogen units were observed, with the highest vibrational frequency being 2385 cm-1, in-line with the previously computed value of 2400 cm-1 for Li2MgH16 at 300 GPa 43. The much higher of Li2MgH16 at 300 GPa compared to that of Li2CaH16 at 350 GPa (3.35 vs. 2.58) is likely in part due to the lower pressure at which it could be stabilized, as was shown to decrease with increasing pressure.43 The larger and lighter mass of Mg as compared to Ca are both factors that result in the higher of Li2MgH16. Another compound that resembled Li2CaH16, LiNa3H23with a type-I SiCL structure, possessed a very similar , and therefore a very similar of 310 K at 350 GPa was predicted 41.
System | Pressure (GPa) | (K) | () | (K) | ||
Li2CaH16 | 350 | 2.58 | 1123 | 1.23 | 330 | 0.94 |
Li2CaH17 | 300 | 2.56 | 1312 | 0.99 | 370 | 1.22 |
Li2CaH17 | 220 | 3.42 | 922 | 0.84 | 317 | 1.42 |
Li2CaH17 | 160 | 1.72 | 1205 | 0.74 | 205 | 1.24 |
@@
Let us now turn to Li2CaH17, whose EPC was well above the threshold requiring the use of the Eliashberg formalism in the studied pressure range: 2.56 at 300 GPa and 1.72 at 160 GPa. At 300 GPa the predicted , 370 K, was near the boiling point of water on Earth’s surface, rendering this another example of a “hot” superconductor. This critical temperature, while lower than that predicted for Li2MgH16, has greatly improved upon the of CaH6. At this pressure vibrations that fell in the primarily Ca-based low frequency region contributed 15% to the total EPC, whereas Li-dominant vibrations contributed 4%, and primarily H vibrations contributed 81%. The superconducting properties of Li2CaH17 can be compared directly to Li2ScH17, Li2YH17, Li2LaH17, and Li2NaH17, hydride superconductors with similar structures 27, 41 (Table S8). Li2CaH17 and Li2NaH17 not only stand out as being better metals than Li2ScH17, Li2YH17, and Li2LaH17, but also have much larger EPCs, leading to much higher predicted s.
@@
Both the and the decreased with decreasing pressure for Li2CaH17, although was highest at 220 GPa, where was lowest. Analysis of the structure and the ELF (Figure S7) revealed that the hydrogen clathrate is beginning to break apart into molecular units at this pressure, leading to the large increase in caused by the structural instability. However, these molecular units reduce in-line with previously observed trends 22. The formation of molecular units becomes more pronounced as pressure decreases, and by 160 GPa 50% of the hydrogen atoms constitute H2 or H molecules, with the other half of the hydrogens forming 9-membered rings, decreasing and the calculated even further. The for Li2CaH17 at 160 GPa is calculated to be 205 K, surprisingly similar to that of CaH6 ( = 210-215 K, 160-172 GPa) 8, 9, 10. At 160 GPa, the limit of dynamical stability for Li2CaH17 (Figure S14), its superconducting figure of merit 77, 1.24, is markedly similar to that of H3S at 150 GPa, 1.31, and somewhat larger than that of Li2LaH17 at 160 GPa, 0.95.3, 27 The highest phonon modes for Li2CaH17 are 2275 cm-1 at 300 GPa and 2190 cm-1 at 160 GPa. These modes can be compared to the highest phonon modes for atomic metallic hydrogen at 500 GPa being approximately 2600 cm-1 78, with those of LaH10 being 2300 cm-1 at 300 GPa 4, 5 and for CaH6 being near 2000 cm-1 at 150 GPa10.

### v2

@@ 3.2 Ternary Phase Diagrams
The most stable structures found in the CSP searches were employed to construct 3D convex hulls (Figures 1, S1-S3) at 100, 200, and 300 GPa at zero-temperature and within the clamped nuclei approximation. These plots are based upon the formation enthalpy as a function of the stoichiometry. By definition the convex hull connects phases whose enthalpies are lower than that of any other phase or linear combination of phases at a particular composition. Thus, the phases that lie on the hull are thermodynamically stable, whereas those that lie above it may be important local minima provided they are close to the hull, are dynamically stable, and are protected by sufficiently high barriers. Therefore, it is often useful to plot the distance of a particular structure from the hull, to determine if it could be an important metastable phase. Because the compositions considered all possessed a 2:1 lithium to calcium ratio while the hydrogen content varied, 2D cuts of the generated convex hulls, corresponding to a plane of increasing H content from Li2Ca at 100 GPa, or 2 Li + Ca at 200 and 300 GPa were also plotted (Figure S5, S6). At 300 GPa a Li2CaH17 phase was predicted to be thermodynamically stable, while the structure closest to, but not on, the hull, Li2CaH16, lay 7 meV/atom above it (Table S2 and Figure 2). Notably, as discussed above, both of these possessed hydrogenic clathrate cages.
@@
The zero-point-energy (ZPE) is often important in determining relative energy orderings for systems containing light elements, with the greatest potential influence on H-bearing phases, as they can have the highest frequency vibrations. For a given pressure the ZPE is affected by symmetry, with lower symmetry systems generally having larger ZPEs, and the nature of the hydrogenic motifs present within the crystal, with dihydrogen molecules having a larger ZPE than hydrogens within extended lattices such as clathrate cages. Indeed, it has been shown that the ZPE can affect the identity of the on-hull structures in high-pressure polyhydrides containing electropositive metal atoms 71. Obtaining the ZPE necessitates performing phonon calculations, which allows for confirmation of dynamic stability, indicated by a lack of imaginary modes.
@@
Phonon calculations, performed in the harmonic approximation, suggested that Li2CaH17, with a hydrogenic clathrate lattice, and Li2CaHn, = (10,12,15,18-20), which did not feature clathrate structures, were dynamically stable at 300 GPa. Li2CaH16, on the other hand, possessed two modes with imaginary frequencies. These imaginary frequencies are related to the acoustic modes, with the majority of the atomic displacement being due to hydrogen. These frequencies were found to become real and positive by 350 GPa (Figure S21), indicating stabilization due to pressure. A first-principles molecular dynamics simulation at 330 K, however, showed that the hydrogenic lattice remained intact, with no molecular H units forming. This suggests one or more of the following: the phase is dynamically stable only at finite temperature; the system may be stabilized by the inclusion of anharmonicity or quantum effects; or the fineness of the grid needed to obtain real phonons is prohibitively computationally expensive. Due to the large size of the unit cell of Li2CaH16, and the associated computational expense, further calculations investigating this discrepancy and determining the lower boundary of dynamic stability were not performed, and the ZPE for this phase was estimated neglecting the imaginary modes. Figure 2(b) illustrates that inclusion of the ZPE lowered the enthalpy of Li2CaH16 so that it joined Li2CaH17 on the 300 GPa hull, but the remaining non-clathrate structures became destabilized, lying at least 40 meV/atom above the hull.
@@
Decreasing pressure moves all of the ternary hydrides progressively further from the hull. Li2CaH16 shifts up to be 16 and 35 meV/atom above it, while Li2CaH17 is calculated to be 9 and 36 meV/atom shy of thermodynamic stability at 200 and 100 GPa, respectively (Figures S1-S6). However, these compounds may still be promising at sub-300 GPa pressures, despite not laying on the hull, provided they are dynamically and thermally stable. Analysis of the zero-temperature DFT-computed enthalpies of the compounds contained in the Materials Project Database led to the estimate that 50% of known inorganic crystalline materials are metastable, with a median metastability energy of 15 meV/atom and a 90 percentile of 67 meV/atom 79. For the specific case of high-pressure hydrides, whose synthesis at high-pressures in diamond anvil cells is often catalyzed by non-equilibrium processes such as thousands-of-Kelvin degree laser heating, numerous examples of synthesized DFT-metastable compounds are known. Some examples include Ca2H577, which lays 20 meV/atom above the hull at 25 GPa, and various PHn phases, which lay 30 meV/atom above the hull within the pressure range 80-225 GPa 80, 81, 82, 83.
@@ 3.3 Hydrogen Clathrate Structures
Plots of the enthalpy versus pressure of the low lying structures found in our CSP searches (Figures S7, S8) showed that Li2CaH16 became preferred (over other phases with this stoichiometry) above 270 GPa, while for Li2CaH17 the structure was the most stable above 140 GPa. From the other candidates found in our EA searches, these had the lowest enthalpies up to a maximum investigated pressure of 500 GPa. The propensity for high in compounds with hydrogenic clathrate lattices, the observation that Li2CaH16 and Li2CaH17 both possessed such lattices with a high degree of symmetry, and the finding that both lay on the 300 GPa convex hull upon inclusion of the ZPE, prompted us to examine them in more detail.
@@
Li2CaH16 is isotypic with previously predicted Li2MgH16 and Li2NaH16 42, 44. In it, each Ca atom lies in a cage surrounded by four hexagons and twelve pentagons, made of 28 hydrogen atoms, and each Li atom is enclathrated in an 18 atom hydrogen cage with four hexagons and five pentagons. The size of the cages around each metal atom are commensurate with their radii, with Ca nestled in the larger cage and lithium sitting in the smaller cage, as illustrated in Figure 3. At 350 GPa these polygons are highly regular, with H-H bond lengths of 1.05, 1.08, and 1.12 Å. This symmetry lattice is markedly similar to the type-II SiCL structure, differing by a single Wyckoff position. While the clearest hydrogen motifs in Li2CaH16 are the H18 and H28 cages, we can also view the structure as possessing eight H4 tetrahedra in the conventional cell ( = 8), corresponding to one H4 tetrahedra per formula unit. These H4 tetrahedra are centered around the unfilled Wyckoff position, a clathrate atom at the (8a) 1/8, 1/8, 1/8 site, that if filled would make them isotypic to the type-II SiCL structure.
@@
In Li2CaH17, also shown in Figure 3, each Ca atom is surrounded by four hexagons and twelve pentagons, and each Li atom is once again found in the smaller cage, surrounded by eleven pentagons. The additional hydrogen atom in Li2CaH17 renders the phase nearly isostructural with the type-II SiCLs. In comparison with Li2CaH16, Li2CaH17 possesses H4 tetrahedra “stuffed” with the extra hydrogen atom, which assists in completing the cage around the Li ions. The stuffed hydrogen atom may stabilize this structure by balancing the chemical pressure interactions within the lattice, as has been observed in the H-stuffed H4 tetrahedra of LaH10 70, or by allowing the structure to distort more readily. These distortions can be seen with the reduced symmetry of the faces comprising the polygons that make up the clathrate; each pentagon has H-H bond lengths between 1.08 and 1.14 Å, while each hexagon has H-H bond lengths between 1.02 and 1.14 Å at 300 GPa. Investigation of the Li2CaH17 phase shows that these bonds become less uniform as pressure decreases. The phase was determined to be dynamically stable down to 160 GPa, however the full clathrate structure does not persist to this pressure, due to the H-H bond lengths increasing (e.g. see ELF plot in Figure S20).
@@
Li2CaH17 was not anticipated to be the most stable spacegroup for this stoichiometry. Previous studies have predicted an M3H17 phase, where M is a metal atom, in compounds such as Li2(Sc/La/Y/Na)H17 28, 42. These symmetry M3H17 phases are isostructural with the type-II SiCL, while the Li2CaH17 phase can be seen as a distortion of this structure. The external pressure, internal (or chemical pressure) and the number of valence electrons that are transferred from the metal atoms to the hydrogen cage, and potential metal-hydrogen interactions may be important factors driving this distortion. The external pressure most directly affects the (pressure-volume) term to the enthalpy, often favoring the formation of closed packed structures at higher pressures. The chemical pressure is related to the ionic radius of the metal atom and how it fits into the cage, with ions that are too-small or too-large resulting in distorted lattices 70. The number of valence electrons that are donated from the electropositive element to the hydrogenic lattice can influence the types of hydrogenic motifs that are present 11, and the resulting H-H distances, which are also influenced by interactions such as Ca d H2 backbonding 77. To confirm that the phase is indeed favored over , the cubic structure was built and optimized. Comparison of the 300 GPa enthalpies revealed that the phase possessed a lower internal energy, , but a higher term. The bonds in the phase are more distorted than the highly ordered phase, and these distortions likely contribute to the enthalpic favorability of the structure for the metal atoms studied here. Furthermore, the phase of Li2CaH17 was determined to be both dynamically unstable at 300 GPa, and enthalpically disfavored over the pressure range of 100-500 GPa, so it was therefore not considered for additional calculations.
@@
A similar phenomenon of competing phases has recently been investigated in the LaH10 system 84. While DFT calculations with classical nuclei and neglecting anharmonic effects predicted that an phase becomes unstable with respect to distorted phases including rhombohedral LaH10 below 230 GPa, the inclusion of quantum fluctuations resulted in the dynamic stability of the phase across a wide pressure regime. Moreover, the competing phases all converged to LaH10 when quantum fluctuations were applied in the structural relaxation. While it is possible that quantum fluctuations could stabilize Li2CaH17 over the phase, these calculations were not performed for several reasons. First, an increased measure of non-uniformity among the H-H bonds that make up the clathrate faces is observed for FPMD simulations at 370 K, which is in alignment with the phase, but not the highly symmetric phase. Since the thermal energy provided by the thermostat is far in excess of the stabilization energy due to quantum fluctuations in hydrides, this distortion would be expected if the cubic structure was preferred. Moreover, at no point over the pressure regime did the phase become thermodynamically preferred over the phase, in contrast to the behavior observed for LaH10. Finally, the computational expense required was seen as overly prohibitive, though recently developed machine-learning approaches may make these types of calculations possible in the future 85. Compared to molecular H-H bonds at the same pressure, the H-H bond lengths in both Li2CaH17 and Li2CaH16 are longer (Table 2). Bader analysis showed that the caged Li and Ca atoms donate charge to the nearby hydrogen atoms, which may covalently bond with one another. For Li2CaH17 at 160 GPa, each H atom gains on average 0.16, with Ca losing 1.05 and Li donating a smaller amount of charge, 0.84. When the pressure is increased to 300 GPa, each H atom in Li2CaH17 gains on average 0.15, while Ca loses 0.89 and Li donates 0.80. Thus, increasing pressure appears to even out the amount of charge transferred from both the formally divalent and monovalent metal atoms, perhaps because a shorter distance between the hydrogen atom and calcium can enhance the amount of backdonation between the occupied antibonding hydrogen-lattice states and the Ca d orbitals, as previously found within compressed CaH4 77 and Ca-S-H ternary hydrides 86. The Bader charges calculated for Li2CaH16 at 350 GPa are similar, with H gaining 0.15 on average, and Ca and Li donating 0.90 and 0.78, respectively.
@@ 3.4 Electronic Structure and Superconductivity
Li2CaH16 and Li2CaH17 are calculated to have a high (Figure 4), primarily of hydrogen character, boding well for superconductivity. The contribution of the calcium states to is intermediate, while the contribution of the lithium states is minimal. Therefore, the DOS plots are consistent with a Li+ cation, and a not-fully-ionized calcium, resulting from the back-donation of charge from the hydrogenic lattice to the Ca d orbitals, as supported by the large amount of Ca d states at . In Li2CaH16 at 350 GPa and Li2CaH17 at 300 GPa lies just below a peak in the DOS, suggesting that further electron doping (e.g. by replacing Li via a divalent metal atom with a similar ionic radius, or Ca by a trivalent metal atom) could potentially enhance . Such peaks in the DOS have previously been reported in symmetry Li2NaH17 (just above ) and Li2YH17 (just below ) at 300 GPa 28, 42. Assuming a full transfer of electrons from the electropositive element to the hydrogenic lattice yields three-donated-electrons in Li2NaH17 and four-donated-electrons in Li2CaH17. In both of these systems sits below the peak in the DOS, so more electrons must be added. However, in Li2YH17, with five-donated-electrons, lies just above the peak. Thus, the optimal number of electron required to maximize is between four and five. We find that adding an additional 0.38 per FU (formula unit) (4 /10 FU) to Li2CaH16 at 350 GPa and an additional 0.40 per FU (4 / 10 FU) to Li2CaH17 at 300 GPa would place on that peak. Therefore, a quaternary compound with an approximate Li2Ca0.6M0.4H16/17 stoichiometry, where M is a trivalent metal ion such as Sc, Y, or La may potentially result in a higher than those calculated here, provided the structure is dynamically stable.
The descriptors for high- superconductivity – and the logarithmic average phonon frequency, – were computed for these two phases at pressures where they were dynamically stable (Table 3). The parameter provides the strength of the electron-phonon coupling in conventional superconductors and falls in the exponential in the semi-empirical Allen-Dynes modified McMillan equation often used to estimate (Equation 1). The parameter lies in the pre-factor of Equation 1. A short overview of the various methods that are commonly used to estimate can be found in Ref. 23.
@@
For Li2CaH16, imaginary modes were absent only at 350 GPa, where this phase possessed the largest that we computed. The large of 2.58 necessitated that the Eliashberg equations be solved numerically, and using a value of 0.1 for the Coulomb shielding parameter, , we obtained a of 330 K for Li2CaH16 at this pressure. Though this predicted is over 100 K lower than that computed for the isotypic Li2MgH16 at 250 GPa 44, it surpasses room temperature, rendering Li2CaH16 a “hot” superconductor. Its phonon band structure, phonon projected densities of states and Eliashberg spectral function are shown in Figure 5(a). Because of the mass differences between the three atom types, the phonons can be well separated into regions where the contributions from a particular atom are dominant. Therefore, the total EPC could be partitioned into contributions originating from vibrations in the low-frequency regime, projected mostly onto the calcium atoms (15%), followed by a lithium-dominant region (9%), then by the mainly hydrogenic modes above 750 cm-1 (76%). No high-frequency modes indicative of the presence of molecular hydrogen units were observed, with the highest vibrational frequency being 2385 cm-1, in-line with the previously computed value of 2400 cm-1 for Li2MgH16 at 300 GPa 44. The much higher of Li2MgH16 at 300 GPa compared to that of Li2CaH16 at 350 GPa (3.35 vs. 2.58) is likely in part due to the lower pressure at which it could be stabilized, as was shown to decrease with increasing pressure.44 The larger and lighter mass of Mg as compared to Ca are both factors that result in the higher of Li2MgH16. Another compound that resembled Li2CaH16, LiNa3H23with a type-I SiCL structure, possessed a very similar , and therefore a very similar of 310 K at 350 GPa was predicted 42.
System | Pressure (GPa) | (K) | () | (K) | (K) | ||
Li2CaH16 | 350 | 2.58 | 1123 | 1.23 | 272 | 330 | 0.94 |
Li2CaH17 | 300 | 2.56 | 1314 | 0.99 | 283 | 370 | 1.22 |
Li2CaH17 | 220 | 3.42 | 922 | 0.84 | 286 | 317 | 1.42 |
Li2CaH17 | 160 | 1.72 | 1194 | 0.74 | 179 | 205 | 1.24 |
@@
Let us now turn to Li2CaH17, whose EPC was well above the threshold requiring the use of the Eliashberg formalism in the studied pressure range: 2.56 at 300 GPa and 1.72 at 160 GPa. At 300 GPa the predicted , 370 K, was near the boiling point of water on Earth’s surface, rendering this another example of a “hot” superconductor. This critical temperature, while lower than that predicted for Li2MgH16, has greatly improved upon the of CaH6. At this pressure vibrations that fell in the primarily Ca-based low frequency region contributed 15% to the total EPC, whereas Li-dominant vibrations contributed 4%, and primarily H vibrations contributed 81%. The superconducting properties of Li2CaH17 can be compared directly to Li2ScH17, Li2YH17, Li2LaH17, and Li2NaH17, hydride superconductors with similar structures 28, 42 (Table S8). Li2CaH17 and Li2NaH17 not only stand out as being better metals than Li2ScH17, Li2YH17, and Li2LaH17, but also have much larger EPCs, leading to much higher predicted s. Calculations performed with the Allen-Dynes modified McMillan equation for strong coupling (Equation 1) yielded lower estimates of the .
@@
Both the and the decreased with decreasing pressure for Li2CaH17, although was highest at 220 GPa, where was lowest. Analysis of the structure and the ELF (Figure S19) revealed that the hydrogen clathrate is beginning to break apart into molecular units at this pressure, leading to the large increase in caused by the structural instability. However, these molecular units reduce in-line with previously observed trends 23. The formation of molecular units becomes more pronounced as pressure decreases, and by 160 GPa 50% of the hydrogen atoms constitute H2 or H molecules, with the other half of the hydrogens forming 9-membered rings, decreasing and the calculated even further. The for Li2CaH17 at 160 GPa is calculated to be 205 K, surprisingly similar to that of CaH6 ( = 210-215 K, 160-172 GPa) 9, 10, 11. At 160 GPa, the limit of dynamical stability for Li2CaH17 (Figure S23), its superconducting figure of merit 87, 1.24, is markedly similar to that of H3S at 150 GPa, 1.31, and somewhat larger than that of Li2LaH17 at 160 GPa, 0.95.4, 28 The highest phonon modes for Li2CaH17 are 2275 cm-1 at 300 GPa and 2190 cm-1 at 160 GPa. These modes can be compared to the highest phonon modes for atomic metallic hydrogen at 500 GPa being approximately 2600 cm-1 88, with those of LaH10 being 2300 cm-1 at 300 GPa 5, 6 and for CaH6 being near 2000 cm-1 at 150 GPa11.